Taylor Corporation reports inventory and cost of goods sold based on calculations from a LIFO periodic inventory system. The company’s records under this system reveal the following inventory layers at the beginning of 2021 (listed in chronological order of acquisition):

18,500 units @ $20 $ 370,000
23,500 units @ $25 587,500
Beginning inventory $ 957,500


During 2021, 47,000 units were purchased for $30 per unit. Due to unexpected demand for the company's product, 2021 sales totaled 58,000 units at various prices, leaving 31,000 units in ending inventory.

Required:
1. Calculate the amount to report for cost of goods sold for 2021.
2. Determine the amount of LIFO liquidation profit that the company must report in a disclosure note to its 2021 financial statements. Assume an income tax rate of 25%.
3. If the company decided to purchase an additional 11,000 units at $30 per unit at the end of the year, how much income tax currently payable would be saved?
Answer #1

Given Information

Units sold during 2021 = 58000 units

Purchases during 2021 = 47000 units

Opening Inventory = 42000 units

1. Calculation of amount of cost of goods sold for 2021 under LIFO Method

Cost Year Quantity sold Quantity Remaining Cost/unit COGS
2021 47000 0 $ 30 $ 1410000
2020 11000

12500 units @ 25

18500 units @ 20

$ 25 $ 275000
Cost of goods sold $ 1685000

Cost of Inventory under LIFO Method = $ 682500

(18500 * 20) + (12500 * 25)


2)

LIFO liquidation refers to the practice of selling or issuing of older merchandise stock or materials in a company’s inventory. It is done by companies that are using the LIFO (last in, first out) inventory valuation method. The liquidation occurs when a company using LIFO sells more goods or issues more old stock than it buys.

Calculation of LIFO Liquidation profit

Given Current year price = 30

Total units sold = 58000

Total cost of Goods sold based on current year price = 58000 * 30 = $ 1740000

Cost of goods sold based on LIFO method = $ 1685000

LIFO liquidation profit = (1740000 - 1685000)

= $ 55000

As tax rate given = 25%

Because of LIFO liquidation we are going to get $ 55000 more profit

So tax payable on LIFO liquidation profit = $ 55000 * 25%

= $ 13750

3) If we had made a purchase of 11000 units at the end of the year 2021

Total purchases made during the year = ( 47000+ 11000) = 58000

Therefore, total sales will be done with current year purchases then there will be no LIFO liquidation

so we won't get $ 55000 more profit so there is no need to pay more tax

Hence , tax payable saved = $ 13750